In a 3 digit number unit digit is one more than the hundred digit are in a ratio 1:2:3 .if the diff of original number and the number obtained by reversing the digit is 594 find the number
Answer by ankor@dixie-net.com(22740)

let a = the hundreds digit
let b = the tens
let c = the units
In a 3 digit number, unit digit is more than the hundred digit,
c > a
:
are in a ratio 1:2:3; therefore:
b = 2a
c = 3a
:
if the diff of original number and the number obtained by reversing the digit is 594,
Reverse number - original number 594
(100c + 10b + a) - (100a + 10b + c) = 594
100c + 10b + a - 100a - 10b - c = 594
Combine like terms
99c - 99a = 594
simplify, divide by 99
c - a = 6
we know c = 3a
3a - a = 6
2a = 6
a = 3
then
b = 2(3) = 6
c = 3(30 = 9
:
find the number: 369 is the original number
:
:
See if that checks out
963 - 369 = 594